OpenStudy (naveenbhatia1312):

Find the exact value by using a half-angle identity. cos(-pi/8)

OpenStudy (naveenbhatia1312):

@misty1212

OpenStudy (misty1212):

hi!!

OpenStudy (misty1212):

\[-\frac{\pi}{8}\] is half of \(-\frac{\pi}{4}\) use the half angle formula with that one

OpenStudy (dayakar):

\[\cos (-\theta) = \cos \ theta]

OpenStudy (dayakar):

\[\cos (-\theta) = \cos \theta\]

OpenStudy (dayakar):

\[\cos (-\frac{ \pi }{ 8 }) = \cos \frac{ \pi }{ 8 }\]

OpenStudy (misty1212):

\[\cos(-\frac{\pi}{8})=\sqrt{\frac{1+\cos(-\frac{\pi}{4})}{2}}\]

OpenStudy (misty1212):

or as @dayakar pointed out, you can use \[\cos(\frac{\pi}{8})=\sqrt{\frac{1+\cos(\frac{\pi}{4})}{2}}\]

OpenStudy (misty1212):

you know \(\cos(\frac{\pi}{4})\) ?

OpenStudy (naveenbhatia1312):

rad 2/2

OpenStudy (dayakar):

\[\cos \frac{ \pi }{ 8 } = \cos \frac{ 180 }{ 8 }\] cos 45

OpenStudy (dayakar):

what is cos 45 value

OpenStudy (naveenbhatia1312):

sq root 2/2

OpenStudy (dayakar):

no check again

OpenStudy (naveenbhatia1312):

.707

OpenStudy (dayakar):

what is that

OpenStudy (naveenbhatia1312):

?

OpenStudy (dayakar):

\[\cos 45 =\frac{ 1 }{ \sqrt{2} }\]

OpenStudy (misty1212):

yes you are right, it is \(\frac{\sqrt2}{2}\)

OpenStudy (misty1212):

\[\cos(\frac{\pi}{8})=\sqrt{\frac{1+\frac{\sqrt2}{2}}{2}}\]

OpenStudy (misty1212):

simplify the compound fraction by multiplying top and bottom by 2 inside the radical

OpenStudy (naveenbhatia1312):

sq root 2/2

OpenStudy (naveenbhatia1312):

\[1/2\sqrt{1+\sqrt{2}}\] would this be the answer?

OpenStudy (naveenbhatia1312):

@misty1212